# Critical Learning Periods: Leveraging Early Training Dynamics for Efficient Data Pruning

### v1

@@ 5.2 Translation quality estimators are less effective for German
We find translation quality estimators to be less effective in pruning German than Swahili. For instance, Figure 2(a) indicates that random data pruning outperforms translation quality estimators at 90% and 70% pruning level for both test sets. However, with Swahili, we observe a different behaviour, where LABSE outperforms random pruning on all pruning levels, as depicted in Figure 2(b). Interestingly, pruning Swahili data with LaBSE and LASER yields superior performance compared to training on the full data set, particularly for the 50% pruning level for the FLORES test set. Further, translation quality estimators consistently outperform random pruning in Swahili and do not sacrifice performance even at 90% pruning level with exception of LASER on the MAFAND test set. We believe that these differences in behavior can be attributed to varying data quality between the two languages. The German dataset contain sentences with high variance in sentence length and included sentences with foreign language. Conversely, the Swahili dataset consists of processed automatically aligned sentences, where there is a low sentence length variance and the sentences do not contain any foreign language. See Figure 7 and Figure 8 in the Appendix. Therefore, selecting the best sentences ensures inclusion of high quality translation pairs.
@@
Figure 2: Performance of the translation quality metrics for both German and Swahili. Quality estimation metrics perform better than random in automatically-aligned setting (Swahili) than in the high-quality setting (German).
@@ 5.3 LLM perplexity techniques are competitive for Swahili but sub-par for German
Figure 3 illustrates that utilising LLMs for pruning provides competitive results for Swahili but shows sub-par performance for German. We further observe that LLM perplexity technique across different BLOOM model sizes performs superior compared to random pruning across pruning levels for Swahili (see Figure 3(b)). Among the various BLOOM model sizes, the smallest model, BLOOM 560M consistently outperforms the others, followed by the BLOOM 1B model. It is important to note that LLM techniques perform worse than random for German. This suggests that in cases where limitations in data and compute double bind, utilising pretrained LLMs can save on the limited resources available.
@@
Figure 3: Performance of the LLM perplexity metrics for both German and Swahili. For German, the LLM perplexity metrics perform worse than random whereas for Swahili, the LLM perplexity metrics are approximately the same as random.
@@ 5.4 What pruning techniques work best?
Figure 4 shows the comparison of random, LaBSE and CAT-DIFF pruning techniques for German, French and Swahili. It is evident that CAT-DIFF, outperforms other methods for German, whereas LaBSE demonstrate superior performance for Swahili. Specifically, CAT-DIFF(1,5) surpasses the other techniques for German at each pruning level and retains 99% of the full data’s performance even at 50% prune level (refer to Figure 4(a)). We observe similar trends for French in Figure 4(b) with CAT-DIFF(1,5) also generally outperforms LaBSE and random selection, except for FLORES at the 50% prune level. Conversely, for Swahili, LaBSE is consistently superior to other pruning techniques, followed by CAT-DIFF(1,5) (see Figure 4(c)). We also find that pruning using LaBSE for Swahili yields competitive performance as training on the full dataset.
We also report the statistical significance results with paired bootstrap resampling(Koehn, 2004) for SacreBLEU and COMET in the Appendix, see Table 5, Table 4 and Table 6.
@@ 6 Analysis of Sentence Selection for German, French, and Swahili
Figure 6 shows that CAT-DIFF(1,5) selects sentences that have more unique and rare words as opposed to German that is less lexically diverse and thus results in lower mean frequency. Similarly, CAT-DIFF(1,5) also selects more unique and rare words for Swahili. Although LaBSE performs relatively better in selecting more unique and rare words for French, it still falls short compared selection of words using CAT-DIFF(1,5) for German. This implies that CAT-DIFF(1,5) selection is far more complex than selecting the longest sentences. We leave deeper investigation of sentences selected using various techniques to future studies.
Figure 5: Target sentence length vs BLEU score for all languages and techniques. We see that longer sentences correlate on average with better performance. Notably, pruning techniques like CAT-DIFF(1,5), which favour longer sentences, achieve higher BLEU scores.
@@ 7 Conclusions
Identifying optimal subsets of a given dataset for training a machine translation model is a crucial problem that enables data efficiency and saves on compute. In this paper, we present Checkpoints Across Time (CAT) methods as efficient techniques to prune datasets. CAT uses the change in perplexity scores of individual examples over the first few training epochs to find good subsets. Using 3.8M sentences for English-German, English-French and English-Swahili, we show that CAT techniques offer highly computationally efficient pruning that achieves above 75% of full performance even when 90% pruning is applied to all target languages. We also show that LaBSE, when used to select training examples from an automatically aligned dataset provides strong signals on examples crucial for model generalisation as the examples selected perform on par or surpass training on the full dataset. Interestingly, we find that utilizing LLMs for pruning does not offer benefits in selecting data for German but does for Swahili.
@@ Appendix A Dataset Analysis
Figure 8: English, en, sentences Language ID vs Target, tgt,(German, Swahili, French) sentence length. Swahili set does not contain any foriegn languages while German and French seem to contain other languages.
@@ Appendix B Other evaluation metrics and significance tests
We ran additional experiments to calculate chrF++ and COMET scores and report the statistical significance results with paired bootstrap resampling (Koehn, 2004) using sacreBLEU and COMET libraries with default configurations for the respective metrics below for all our experiments with on the 3.8M data scale. We find that CAT-DIFF demonstrates competitive or superior performance compared to other methods on en-de and en-fr pairs across most test sets. Bolded numbers in Table 4, Table 5 and Table 6 show that with 1000 resamples, CAT-DIFF outperforms the other technique 95% of the time thus CAT-DIFF is statistically significance at 95% confidence level. These results highlight the effectiveness of our proposed method.
@@
Table 4: Significance test for Swahili with the null hypothesis: mean translation score of CAT-DIFF is equal to the mean of the other technique. A bolded entry signifies that CAT-DIFF is significant better (p-value $<0.05$) than the method used for the corresponding value. For each dataset, we show results for different prune levels (90%, 70%, 50%) and metrics (BLEU, chrF++, COMET). Here "Full" means the value obtained when using all the training data.
@@
Table 5: Significance test for German with the null hypothesis: mean translation score of CAT-DIFF is equal to the mean of the other technique. A bolded entry signifies that CAT-DIFF is significant better (p-value $<0.05$) than the method used for the corresponding value. For each dataset, we show results for different prune levels (90%, 70%, 50%) and metrics (BLEU, chrF++, COMET). Here "Full" means the value obtained when using all the training data.
@@
Table 6: Significance test for French with the null hypothesis: mean translation score of CAT-DIFF is equal to the mean of the other technique. A bolded entry signifies that CAT-DIFF is significant better (p-value $<0.05$) than the method used for the corresponding value. For each dataset, we show results for different prune levels (90%, 70%, 50%) and metrics (BLEU, chrF++, COMET). Here "Full" means the value obtained when using all the training data.

### v2

@@ 5.2 Translation quality estimators are less effective for German
We find translation quality estimators to be less effective in pruning German than Swahili. For instance, Figure 2(a) indicates that random data pruning outperforms translation quality estimators at 90% and 70% pruning levels for both test sets. However, with Swahili, we observe a different behavior, where LABSE outperforms random pruning on all pruning levels, as depicted in Figure 2(b). Interestingly, pruning Swahili data with LaBSE and LASER yields superior performance compared to training on the full data set, particularly for the 50% pruning level for the FLORES test set. Further, translation quality estimators consistently outperform random pruning in Swahili and do not sacrifice performance even at 90% pruning level with the exception of LASER on the MAFAND test set. We believe that these differences in behavior can be attributed to varying data quality between the two languages. The German dataset contains sentences with high variance in sentence length and included sentences with foreign language. Conversely, the Swahili dataset consists of processed automatically aligned sentences, where there is a low sentence length variance and the sentences do not contain any foreign language. See Figure 7 and Figure 8 in the Appendix. Therefore, selecting the best sentences ensures the inclusion of high-quality translation pairs.
@@
Figure 2: Performance of the translation quality metrics for both German and Swahili. Quality estimation metrics perform better than random in an automatically aligned setting (Swahili) than in the high-quality setting (German).
@@ 5.3 LLM perplexity techniques are competitive for Swahili but sub-par for German
Figure 3 illustrates that utilizing LLMs for pruning provides competitive results for Swahili but shows sub-par performance for German. We further observe that the LLM perplexity technique across different BLOOM model sizes performs superior compared to random pruning across pruning levels for Swahili (see Figure 3(b)). Among the various BLOOM model sizes, the smallest model, BLOOM 560M consistently outperforms the others, followed by the BLOOM 1B model. It is important to note that LLM techniques perform worse than random for German. This suggests that in cases where limitations in data and compute double bind, utilizing pretrained LLMs can save on the limited resources available.
@@
Figure 3: Performance of the LLM perplexity metrics for both German and Swahili. For German, the LLM perplexity metrics perform worse than random whereas, for Swahili, the LLM perplexity metrics are approximately the same as random.
@@ 5.4 What pruning techniques work best?
Figure 4 shows the comparison of random, LaBSE and CAT-DIFF pruning techniques for German, French and Swahili. It is evident that CAT-DIFF, outperforms other methods for German, whereas LaBSE demonstrates superior performance for Swahili. Specifically, CAT-DIFF(1,5) surpasses the other techniques for German at each pruning level and retains 99% of the full data’s performance even at 50% prune level (refer to Figure 4(a)). We observe similar trends for French in Figure 4(b) with CAT-DIFF(1,5) also generally outperforms LaBSE and random selection, except for FLORES at the 50% prune level. Conversely, for Swahili, LaBSE is consistently superior to other pruning techniques, followed by CAT-DIFF(1,5) (see Figure 4(c)). We also find that pruning using LaBSE for Swahili yields competitive performance as training on the full dataset.
We also report the statistical significance results with paired bootstrap resampling(Koehn, 2004) for SacreBLEU and COMET in the Appendix, see Tables 5, 4 and 6.
@@ 6 Analysis of Sentence Selection for German, French, and Swahili
Figure 6 shows that CAT-DIFF(1,5) selects sentences that have more unique and rare words as opposed to German which is less lexically diverse and thus results in lower mean frequency. Similarly, CAT-DIFF(1,5) also selects more unique and rare words for Swahili. Although LaBSE performs relatively better in selecting more unique and rare words for French, it still falls short compared selection of words using CAT-DIFF(1,5) for German. This implies that CAT-DIFF(1,5) selection is far more complex than selecting the longest sentences. We leave a deeper investigation of sentences selected using various techniques for future studies.
Figure 5: Target sentence length vs BLEU score for all languages and techniques. We see that longer sentences correlate on average with better performance. Notably, pruning techniques like CAT-DIFF(1,5), which favor longer sentences, achieve higher BLEU scores.
@@ 7 Conclusions
Identifying optimal subsets of a given dataset for training a machine translation model is a crucial problem that enables data efficiency and saves on compute. In this paper, we present Checkpoints Across Time (CAT) methods as efficient techniques to prune datasets. CAT uses the change in perplexity scores of individual examples over the first few training epochs to find good subsets. Using 3.8M sentences for English-German, English-French and English-Swahili, we show that CAT techniques offer highly computationally efficient pruning that achieves above 75% of full performance even when 90% pruning is applied to all target languages. We also show that LaBSE, when used to select training examples from an automatically aligned dataset provides strong signals on examples crucial for model generalization as the examples selected perform on par or surpass training on the full dataset. Interestingly, we find that utilizing LLMs for pruning does not offer benefits in selecting data for German but does for Swahili.
@@ Appendix A Dataset Analysis
Figure 8: English sentences Language ID (en) vs Target (German, Swahili, French) sentence length (tgt). Swahili set does not contain any foreign languages while German and French seem to contain other languages.
@@ Appendix B Other evaluation metrics and significance tests
We ran additional experiments to calculate chrF++ and COMET scores and report the statistical significance results with paired bootstrap resampling (Koehn, 2004) using sacreBLEU and COMET libraries with default configurations for the respective metrics below for all our experiments with on the 3.8M data scale. We find that CAT-DIFF demonstrates competitive or superior performance compared to other methods on en-de and en-fr pairs across most test sets. Bolded numbers in Tables 4, 5 and 6 show that with 1000 resamples, CAT-DIFF outperforms the other technique 95% of the time thus CAT-DIFF is statistically significance at 95% confidence level. These results highlight the effectiveness of our proposed method.
@@
Table 4: Significance test for Swahili with the null hypothesis: mean translation score of CAT-DIFF is equal to the mean of the other technique. A bolded entry signifies that CAT-DIFF is significantly better (p-value $<0.05$) than the method used for the corresponding value. For each dataset, we show results for different prune levels (90%, 70%, 50%) and metrics (BLEU, chrF++, COMET). Here "Full" means the value obtained when using all the training data.
@@
Table 5: Significance test for German with the null hypothesis: mean translation score of CAT-DIFF is equal to the mean of the other technique. A bolded entry signifies that CAT-DIFF is significantly better (p-value $<0.05$) than the method used for the corresponding value. For each dataset, we show results for different prune levels (90%, 70%, 50%) and metrics (BLEU, chrF++, COMET). Here "Full" means the value obtained when using all the training data.
@@
Table 6: Significance test for French with the null hypothesis: mean translation score of CAT-DIFF is equal to the mean of the other technique. A bolded entry signifies that CAT-DIFF is significantly better (p-value $<0.05$) than the method used for the corresponding value. For each dataset, we show results for different prune levels (90%, 70%, 50%) and metrics (BLEU, chrF++, COMET). Here "Full" means the value obtained when using all the training data.